A Honda Civic travels in a straight line along a road. Its distance x from a stop sign is given as a function of time t by the equation x(t)= α t^2− β t^3, where α = 1.42 m/s^2 and β = 5.15×10^−2 m/s^3

a)Calculate the average velocity of the car for the time interval t=0 to t1 = 1.97 s

b)Calculate the average velocity of the car for the time interval t=0 to t2 = 4.10 s

c)Calculate the average velocity of the car for the time interval t1 = 1.97 s to t2 = 4.10 s
Answer #1

a)

Average velocity of the car for the time interval t=0 to t1 = 1.97 s

Xt,-1.97s-Xt=0 t1-t (1.42 x 1.972-5.15 × 10-2 1.973)-0 avg 1.97-0

Vavg=2.5975 m/s =2.6 m/s

b)

Average velocity of the car for the time interval t=0 to t2 =4.1 s

(1.42 × 4.12-5.15 × 10-2 × 4.13)-0 Xt,-41s-Xt=0 t2 -t aug

Vavg=4.9563 m/s =4.96 m/s

c)

Average velocity of the car for the time interval t1=1.97 to t2 =4.1 s

Xt2:41s-X,,-1.97 t2 - ti (1.42x 4.12-5.15 × 10-г × 4.1°) _ (( 1.42 x 1.972-5.15 × 10-2 × 1.973)) .1 1.97 augVavg=7.138 m/s =7.14 m/s